# BnSentMix: A Diverse Bengali-English Code-Mixed Dataset for Sentiment Analysis

### v1

@@ Abstract
The widespread availability of code-mixed data can provide valuable insights into low-resource languages like Bengali, which have limited datasets. Sentiment analysis has been a fundamental text classification task across several languages for code-mixed data. However, there has yet to be a large-scale and diverse sentiment analysis dataset on code-mixed Bengali. We address this limitation by introducing BnSentMix, a sentiment analysis dataset on code-mixed Bengali consisting of 20,000 samples with $4$ sentiment labels from Facebook, YouTube, and e-commerce sites. We ensure diversity in data sources to replicate realistic code-mixed scenarios. Additionally, we propose $14$ baseline methods including novel transformer encoders further pre-trained on code-mixed Bengali-English, achieving an overall accuracy of $69.8\%$ and an F1 score of $69.1\%$ on sentiment classification tasks. Detailed analyses reveal variations in performance across different sentiment labels and text types, highlighting areas for future improvement.
BnSentMix: A Diverse Bengali-English Code-Mixed Dataset for Sentiment Analysis
Sadia Alam, Md Farhan Ishmam, Navid Hasin Alvee, Md Shahnewaz Siddique, Md Azam Hossain, Abu Raihan Mostofa Kamal Network and Data Analysis Group, Department of Computer Science and Engineering, Islamic University of Technology, Dhaka, Bangladesh {sadiaalam,farhanishmam,navidhasin,shahnewaz,azam,raihan.kamal}@iut-dhaka.edu
@@ 1 Introduction
Table 1 highlights the limitations of Bengali-English code-mixed sentiment analysis datasets compared to other Indic-English code-mixed datasets. The only available Bengali dataset Mandal et al. (2018) is limited to $5k$ samples, $3$ sentiment labels, a single data source, $5$ baselines, and is not publicly available. The existing language detection tools also have severe limitations in filtering code-mixed Bengali-English. Tools like langdetect [^1] https://pypi.org/project/langdetect/ and Bengali phonetic parser [^2] https://github.com/porimol/bnbphoneticparser designed for general language identification and code-mixed Bengali identification struggled with the spelling nuances of code-mixed text. The limitations necessitated novel code-mixed Bengali detection tools.
Addressing these challenges, we present a novel and extensive Bengali-English code-mixed dataset designed for sentiment analysis. Our dataset is distinguished for its diversity, volume, and public availability. Data has been curated from YouTube, Facebook, and e-commerce platforms to encapsulate a broad spectrum of contexts and topics. The availability of a robust and diverse dataset is a critical step towards developing more effective and inclusive NLP tools, ultimately contributing to the better understanding and processing of code-mixed languages. Our contributions can be summarized as:
  • We present the first large-scale Bengali-English code-mixed dataset with 20,000 samples and four sentiment labels. The dataset will be publicly available to facilitate further research.
  • We describe an automated Bengali-English code-mixed text detection method using a pre-trained language model for efficient data filtering.
  • We establish 14 baselines including 3 novel transformer encoders further pre-trained on code-mixed Bengali-English, achieving state-of-the-art performance on our dataset.
@@ 2 Related Work
2.1 Code-Mixed Tasks and Methods
Code-mixed data can be the source of several text classification tasks Thara and Poornachandran (2018) with sentiment analysis Mahadzir et al. (2021) being one of the most popular ones. Other natural language processing tasks (NLP) on code-mixed data include hate speech detection Sreelakshmi et al. (2020), translation Gautam et al. (2021), part of speech tagging Vyas et al. (2014), emotion classification Ameer et al. (2022), language identification Mandal and Singh (2018), and speech synthesis Sitaram and Black (2016). Researchers also incorporated several techniques in processing code-mixed data including training data augmentation Gupta et al. (2021); Rizvi et al. (2021) and word embeddings specific to code-mixed text Pratapa et al. (2018).
@@ 2.2 Sentiment Analysis
The significance of sentiment analysis has grown substantially with the rise of social media, prompting extensive research on monolingual corpora. Studies explored various languages, including English Hu and Liu (2004); Wiebe et al. (2005); Jiang et al. (2019), Russian Rogers et al. (2018), German Cieliebak et al. (2017), Norwegian Mæhlum et al. (2019), and several Indian languages Agrawal and Awekar (2018); Rani et al. (2020). Multilingual sentiment analysis Dashtipour et al. (2016); Pustulka-Hunt et al. (2018) gained popularity with the recent advancements in multilingual language models Devlin et al. (2019); Conneau et al. (2020). Over the years, a vast number of code-mixing datasets with English have been proposed, as enlisted in Table 1.
@@ 2.3 Code-Mixing in Bengali
Bengali is often code-mixed with English Chanda et al. (2016) and Hindi Raihan et al. (2023). In Bengali-English code-mixing, English tokens are commonly used alongside romanized or transliterated Bengali Shibli et al. (2023). Sentiment analysis on code-mixed Bengali has limited studies, which either use small private datasets Mandal et al. (2018) or are performed in a multilingual setting Patra et al. (2018). Data augmentation techniques have also been explored to enhance code-mixed sentiment analysis datasets in Bengali Tareq et al. (2023). Emotion detection, a task similar to sentiment analysis, has also been studied in the context of code-mixed Bengali Raihan et al. (2024).
@@ 3 BnSentMix Dataset
3.1 Dataset Design
We designed BnSentMix, a large-scale code-mixed Bengali-English dataset with four sentiment labels. First, we ensured that the raw data was collected from multiple data sources, reflecting realistic code-mixed texts commonly found in digital spaces. Second, we labeled the data using four distinct sentiments: the commonly used positive, negative, and neutral sentiments, as well as a mixed sentiment. As illustrated in Fig. 1, the mixed sentiment represents instances where both positive and negative sentiments are conveyed within different parts of the text. We decided to include the mixed label because the associated sentences are frequently observed in everyday texts and cannot be correctly classified under the traditional sentiment labels.
@@ 3.3 Data Cleaning
We discard samples with four words or less and samples containing external URLs. Redundant whitespaces, special characters, and non-ASCII characters including emojis and emoticons are also removed. Consequent sequences of punctuation symbols are reduced to single instances. The English words are downcased unless they appear at the beginning of the sentence. However, we did not correct any form of typing or grammatical errors in our dataset to ensure the trained model is robust for practical scenarios. The data cleaning procedure excluding the preliminary filtering is described in Algorithm 1.
@@ 3.4 Data Filtering
We construct a novel Bengali-English code-mix detection dataset and fine-tune pre-trained language models to filter code-mixed Bengali-English from the data sources. Pre-trained language models excel in text classification tasks and can address the shortcomings of current rule-based language detection tools.
3.4.1 Code-mix Detection Dataset
The fine-tuning dataset for code-mixed Bengali-English detection comprises $3$ data sources. We incorporate the Dakshina dataset Roark et al. (2020) which has a rich collection of Southeast Asian languages, including many Bengali-English code-mixed sentences. Secondly, we utilized a Kaggle English dataset [^5] https://www.kaggle.com/datasets/rtatman/english-word-frequency consisting of a wide range of English words and extended with a third source Mandal and Singh (2018). By integrating these diverse sources, we curated a comprehensive dataset of $100k$ words, ensuring a balanced mix of Bengali, English, and code-mixed Bengali-English text. To maintain the linguistic purity of code-mixed Bengali, we exclude sentences containing words that are not English or Bengali.
3.4.2 Code-mix Detection Results
Model Acc(%) F1 Score
XLM-RoBERTa 89.60 0.8985
BanglishBERT 90.56 0.8961
mBERT 94.56 0.9403
@@ 3.4.2 Code-mix Detection Results
3.4.3 Filtering Pipeline
Following Fig. 2, the code-mixed Bengali-English is cleaned and passed to the data filtering classifier. The classifier detects the code-mixed Bengali samples from the multilingual data source following Algorithm 2. The filtered code-mixed Bengali is manually verified to ensure the highest quality of data. To maintain linguistic purity, we exclude sentences containing words that are not English or Bengali. We discard samples containing hate speech and offensive content. We also discarded samples with a small proportion of code-mixing. Sentences are considered to be code-mixed Bengali-English if they contain at least 30% English words. We obtain 21,587 data samples after data filtering.
@@ 3.4.3 Filtering Pipeline
Unique Sentence Count 21873
@@
3.5 Data Annotation
Each sample in our dataset has been annotated twice by two different annotators to ensure generalized sentiment is conveyed. To perform data annotation, we recruited $64$ annotators who had been provided hourly monetary compensation. The data annotators have at least a high-school degree (equivalent to Grade 12 education) and are familiar with social media and digital spaces. The annotators were asked to re-label the same $250$ samples. We measured the inter-annotator agreement level and found Cohen’s Kappa $\kappa=0.86$, indicating substantial agreement.
Figure 4: Distribution of sentiment labels in the BnSentMix dataset.
3.6 Annotation Validation
The annotated data has been further verified by a third annotator. In cases where the two independent annotations did not match, the third annotator would break the tie. This additional step ensured that any contentious samples were thoroughly reviewed and the final annotation reflected a consensus among the annotators. Samples labeled unfit for sentiment analysis after data validation have been discarded.
3.7 Dataset Statistics
After a rigorous evaluation process, 20,000 samples were finalized from the 21,587 samples, ensuring high-quality annotations for subsequent analysis and model training. Fig. 4 visualizes the label composition of the annotated dataset. An overview of the key statistics of the annotated dataset is shown in table-3.
3.8 Train-Validation-Test Set Split
To evaluate the performance of our models effectively, we split the dataset into $[70:15:15]$ training, validation, and test splits i.e. 14,000, 3,000, and 3,000 samples respectively.
4 Methodology and Experimental Setup
4.1 Baseline Models
We evaluate $14$ baselines encompassing traditional machine learning models, recurrent neural network variants, and transformer-based pre-trained language models, observed in table 4.
4.2 Fine-tuning Transformer Baselines
While the non-transformer models were trained from scratch, the transformer models leveraged pre-trained weights that were fine-tuned on our dataset. We perform task-specific fine-tuning Howard and Ruder (2018) by initializing the model parameters $\theta_{0}=\theta_{\text{pre}}$, and updating them following the gradient descent equation for $t+1$ iteration:
$\theta_{t+1}=\theta_{t}-\alpha\nabla_{\theta}J(\theta)$ (1)
where $\theta_{i}$ represents the parameters for the $i$ iteration, $\alpha$ is the learning rate and $J(\theta)$ is the derivative of the loss function with respect to the given parameters.
4.3 Further Pre-trained Transformer Models
We enhance the performance of the transformer baselines by incorporating Masked Language Modeling loss Devlin et al. (2019) to further pre-train the models on the unlabeled code-mixed data corpus scraped in sec-3.2. The MLM loss following dynamic masking can be formulated as:
$L_{\text{MLM}}=-\frac{1}{N}\sum_{i=1}^{N}\log P(x_{i}\mid\mathbf{x}_{\text{mask}})$ (2)
where, $x_{i}$ is the token at position $i$ masked in the input sequence, $\mathbf{x}_{\text{mask}}$ denotes the input sequence with dynamically masked tokens, $N$ is the total number of masked tokens. We ensure that samples from the test and validation sets are excluded from the further pre-training corpus. The further pre-trained models are referred to as Code-Mixed Bengali (CMB) transformer models and they follow the naming convention $x$-CMB, where $x$ represents a pre-trained transformer variant.
4.4 Evaluation Metrics
4.4.1 Accuracy
Classification accuracy or accuracy, in short, quantifies the proportion of correctly classified predictions made by the classifier to the total predictions. The metric can be formulated:
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (3)
where, TP, TN, FP, and FN represent the counts of True Positives, True Negatives, False Positives, and False Negatives respectively.
4.4.2 F1 Score
Accuracy can be misleading for imbalanced datasets. Instead, the F1 Score can provide a better estimate, computed as the harmonic mean of precision and recall.
[…]
where,
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (5)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (6)
@@ 4.5 Implementation Details
The models were trained on NVIDIA Tesla P100 GPUs with 16GB of memory. We followed the Huggingface implementation Wolf et al. (2019) for the pre-trained language models. All the models utilized Adam Optimizer Kingma and Ba (2014) with a training batch size of $32$.
The training configuration used most of the default hyperparameters. Logistic Regression, RNN, and LSTM models used the learning rate of $1E{-5}$ while the BERT-family language models used the learning rate of $1.5E{-6}$. The training time for each epoch varied from $8$ to $13$ minutes.
@@
Pre-trained Transformer Models
@@
Code-Mixed Bengali (CMB) Transformer Models
mBERT-CMB 0.727 0.713 0.727 0.720 0.694 0.675 0.695 0.685
BanglaBERT-CMB 0.723 0.669 0.724 0.695 0.699 0.643 0.699 0.670
BERT-CMB 0.730 0.712 0.728 0.720 0.698 0.685 0.698 0.691
@@ 5.1 Performance Evaluation
Table 4 highlights the performance of the $14$ baselines with BERT-CMB being the best-performing model in terms of both accuracy and F1 score. We now analyze the category-wise model performance.
@@ 5.1.3 Transformer-based Models
In contrast, English pre-trained models like BERT exhibit a better understanding of the linguistic intricacies of English words used in code-mixed Bengali, thereby producing better performance than other multilingual and Bengali models. Similarly, the Bengali language models BanglaBERT Bhattacharjee et al. (2022) and BanglishBERT Bhattacharjee et al. (2022) are trained on Bengali and Bengali-English corpora respectively. Code-mixed Bengali uses English tokens and hence, the pre-training on Bengali tokens does not provide any significant advantage. The lighter version of BERT, DistilBERT Sanh et al. (2019) produces comparable but slightly worse results.
5.1.4 Code-Mixed Bengali Models
We selected the three best-performing transformer encoders and further pre-trained them following sec-4.3. The models were selected from each sub-category – mBERT as the best multilingual model, BanglaBERT as the best Bengali model, and BERT as the best English model. We find that further pre-training on the code-mixed data translates to an increase in performance for all three models. The only exception is the mBERT-CMB model showing insignificant improvement in accuracy and F1-score. Similar to the pre-trained baselines, the BERT-CMB is the best-performing model followed by mBERT-CMB and BanglaBERT-CMB. We argue that further pre-training in a larger code-mixed corpus might translate to better performance for multilingual models like mBERT or XLM-RoBERTa.
@@ 5.2 Training Loss Analysis
5.3 Hyperparameter Tuning
The hyperparameters play a crucial role in optimizing the performance of deep learning models. We analyze the role of hyperparameters for the best-performing BERT-CMB model and optimize them to gain superior performance.
5.3.1 Fine-tuning Epochs
Figure 6 shows the effect of fine-tuning epochs on the performance metrics. Both metrics increase up to the $6^{th}$ and stabilize onwards, suggesting keeping the number of fine-tuning epochs between $5$ to $7$.
Figure 6: Epoch-wise training accuracy and F1 score curves of the best performing BERT-CMB model.
5.4 Learning Rate vs. Accuracy Analysis
Figure 7 illustrates the relationship between the learning rate and the accuracy for the BERT-CMB model. While tuning the hyperparameters, we observed the highest accuracy in the validation set at the learning rate of $1.25E{-6}$ which is slightly lower than the default learning rate of $1.5E{-6}$. The curve shows no particular advantage for higher or lower rates.
Figure 7: The effect of learning rate on the validation accuracy of the best performing BERT-CMB model.

### v3

@@ Abstract
The widespread availability of code-mixed data in digital spaces can provide valuable insights into low-resource languages like Bengali, which have limited annotated corpora. Sentiment analysis, a pivotal text classification task, has been explored across multiple languages, yet code-mixed Bengali remains underrepresented with no large-scale, diverse benchmark. Code-mixed text is particularly challenging as it requires the understanding of multiple languages and their interaction in the same text. We address this limitation by introducing BnSentMix, a sentiment analysis dataset on code-mixed Bengali comprising 20,000 samples with $4$ sentiment labels, sourced from Facebook, YouTube, and e-commerce sites. By aggregating multiple sources, we ensure linguistic diversity reflecting realistic code-mixed scenarios. We implement a novel automated text filtering pipeline using fine-tuned language models to detect code-mixed samples and expand code-mixed text corpora. We further propose baselines using machine learning, neural networks, and transformer-based language models. The availability of a diverse dataset is a critical step towards democratizing NLP and ultimately contributing to a better understanding of code-mixed languages.
BnSentMix: A Diverse Bengali-English Code-Mixed Dataset
for Sentiment Analysis
Sadia Alam, Md Farhan Ishmam, Navid Hasin Alvee, Md Shahnewaz Siddique, Abu Raihan Mostofa Kamal, Md Azam Hossain Department of Computer Science and Engineering, Islamic University of Technology {sadiaalam,farhanishmam,navidhasin,shahnewaz,raihan.kamal,azam}@iut-dhaka.edu
@@ 1 Introduction
Table 1 highlights the limitations of Bengali-English code-mixed sentiment analysis datasets compared to other Indic-English code-mixed datasets. The only available Bengali dataset Mandal et al. (2018) is limited to $5k$ samples, $3$ sentiment labels, a single data source, $5$ baselines, and is not publicly available. The existing language detection tools also have severe limitations in filtering code-mixed Bengali-English. Tools like langdetect [^1] https://pypi.org/project/langdetect/ and Bengali phonetic parser [^2] https://github.com/porimol/bnbphoneticparser designed for general language identification and code-mixed Bengali identification struggled with the spelling nuances of code-mixed text.
Addressing these challenges, our contribution can be summarized:
  • We present, BnSentMix, a novel Bengali-English code-mixed dataset comprising 20,000 samples and 4 sentiment labels for sentiment analysis. Data has been curated from YouTube, Facebook, and e-commerce platforms to encapsulate a broad spectrum of contexts and topics.
  • Following the intricacies of code-mixed test, visualized in Fig. 1, we propose a novel automated code-mixed text detection pipeline using fine-tuned language models, reaching an accuracy of 94.56%.
  • We establish 11 baselines including classical machine learning, neural network, and pre-trained transformer-based models, with BERT achieving accuracy and F1 score of 69.5% and 68.8% respectively.
@@ 2 Related Work
2.1 Code-Mixing
Code-mixed data can be the source of several text classification tasks Thara and Poornachandran (2018) with sentiment analysis Mahadzir et al. (2021) being one of the most popular ones. Other natural language processing tasks (NLP) on code-mixed data include hate speech detection Sreelakshmi et al. (2020), translation Gautam et al. (2021), part of speech tagging Vyas et al. (2014), emotion classification Ameer et al. (2022), language identification Mandal and Singh (2018), and speech synthesis Sitaram and Black (2016). Researchers also incorporate training data augmentation Gupta et al. (2021); Rizvi et al. (2021) and code-mix word embeddings Pratapa et al. (2018) to process code-mixed texts.
@@ 2.2 Sentiment Analysis
The significance of sentiment analysis has grown with the rise of social media, prompting extensive research on monolingual corpora. Studies explored various languages, including English Hu and Liu (2004); Wiebe et al. (2005); Jiang et al. (2019), Russian Rogers et al. (2018), German Cieliebak et al. (2017), Norwegian Mæhlum et al. (2019), several Indian languages Agrawal and Awekar (2018); Rani et al. (2020), and Bengali Fahim (2023); Kabir et al. (2023). Multilingual sentiment analysis Dashtipour et al. (2016); Pustulka-Hunt et al. (2018) gained popularity with the recent advancements in multilingual language models Devlin et al. (2019); Conneau et al. (2020).
@@ 2.3 Code-Mixing in Bengali
Bengali is often code-mixed with English Chanda et al. (2016) and Hindi Raihan et al. (2023). In Bengali-English code-mixing, English tokens are commonly used alongside romanized or transliterated Bengali Shibli et al. (2023); Fahim et al. (2024), which is often back-transliterated before processing Haider et al. (2024). Sentiment analysis on code-mixed Bengali has limited studies, either using small private datasets Mandal et al. (2018) or performed in a multilingual setting Patra et al. (2018). Data augmentation techniques have also been explored to enhance code-mixed sentiment analysis datasets in Bengali Tareq et al. (2023). Emotion detection, a task similar to sentiment analysis, has also been studied in the context of code-mixed Bengali Raihan et al. (2024).
@@ 3 BnSentMix Dataset
The BnSentMix data has been collected from multiple data sources to reflect realistic code-mixed texts commonly found in digital spaces. We labeled the data using four distinct sentiments: the commonly used positive, negative, and neutral sentiments, as well as a mixed sentiment. As illustrated in Fig. 1, the mixed sentiment represents instances where both positive and negative sentiments are conveyed within different parts of the text. We decided to include the mixed label because the associated sentences are frequently observed in everyday texts and cannot be correctly classified under the traditional sentiment labels.
@@ 3.2 Data Cleaning
We discard samples with four words or less and samples containing external URLs. Redundant whitespaces, special characters, and non-ASCII characters including emojis and emoticons are also removed. Consequent sequences of punctuation symbols are reduced to single instances. The English words are downcased unless they appear at the beginning of the sentence. However, we did not correct any form of typing or grammatical errors in our dataset to ensure the trained model is robust for practical scenarios. The data cleaning procedure has been formally described in Algo. 1.
@@ 3.3 Data Filtering
We construct a novel Bengali-English code-mix detection dataset and fine-tune pre-trained language models to automatically filter code-mixed Bengali-English. Detecting these texts can pose significant challenges: (i) rule-based methods struggle with intra-word switching (ii) romanized Bengali or English samples may be incorrectly classified as code-mixed text by automated methods, and (iii) samples from a third language often bypass the filtering process. Our approach addresses these challenges by incorporating pre-trained language models, which excel in intricate text detection settings. Algo. 2 outlines the data filtering pipeline.
@@
3.3.1 Code-mix Detection Dataset
The fine-tuning dataset for code-mixed Bengali-English detection comprises $3$ data sources. We incorporate the Dakshina dataset Roark et al. (2020) which has a rich collection of Southeast Asian languages, including many Bengali-English code-mixed sentences. Secondly, we utilized a Kaggle English dataset [^5] https://www.kaggle.com/datasets/rtatman/english-word-frequency consisting of a wide range of English words and extended with a third source Mandal and Singh (2018). By integrating these diverse sources, we curated a comprehensive dataset of $100k$ words, ensuring a balanced mix of Bengali, English, and code-mixed Bengali-English words. To maintain the linguistic purity of code-mixed Bengali-English, we exclude sentences containing words that are neither English nor Bengali, e.g. Hindi words.
3.3.2 Code-mix Detection Results
Model Acc(%) F1 Score
XLM-RoBERTa 89.60 0.8985
BanglishBERT 90.56 0.8961
mBERT 94.56 0.9403
Figure 4: Distribution of sentiment labels in the BnSentMix dataset.
3.4 Data Annotation
Each sample in our dataset has been annotated twice by two different annotators to ensure generalized sentiment is conveyed. In cases where the two independent annotations did not match, a third annotator would break the tie. To perform data annotation, we recruited $64$ annotators who had been provided hourly monetary compensation. The data annotators have at least a high-school degree (equivalent to Grade 12 education) and are familiar with social media and digital spaces. The annotators were asked to re-label the same $250$ samples to measure inter-annotator agreement. We measured the agreement score using Cohen’s Kappa $\kappa=0.86$, indicating substantial agreement.
3.5 Dataset Statistics
Fig. 4 visualizes the label composition of the annotated dataset. An overview of the key statistics of the annotated dataset is shown in table-3. We split the dataset into $[70:15:15]$ training, validation, and test splits i.e. 14,000, 3,000, and 3,000 samples respectively.
4 Methodology and Experimental Setup
4.1 Baseline Models
We evaluate $11$ baselines encompassing traditional machine learning models, recurrent neural network variants, and transformer-based pre-trained language models, observed in table 4. All the pre-trained models were fine-tuned on our dataset.
4.2 Evaluation Metrics
@@
Unique Sentence Count 20000
@@ 4.3 Implementation Details
The models were trained on NVIDIA Tesla P100 GPUs with 16GB of memory. We followed the Huggingface implementation Wolf et al. (2019) for the pre-trained language models. All the models utilized Adam Optimizer Kingma and Ba (2014) with a training batch size of $32$. The training configuration used most of the default hyperparameters. Logistic Regression, RNN, and LSTM models used the learning rate of $1E{-5}$ while the BERT-family language models used the learning rate of $1.5E{-6}$. The training time for each epoch varied from $8$ to $13$ minutes.
@@
Multilingual Language Models
@@ 5.1 Performance Evaluation
Table 4 highlights the performance of the $11$ baselines with BERT achieving the best performance in terms of both accuracy and F1 score. We now analyze the category-wise model performance.
@@ 5.1.3 Transformer-based Models
In contrast, English pre-trained models like BERT exhibit better understanding of the linguistic intricacies of English words used in code-mixed Bengali, thereby producing better performance than other multilingual and Bengali models. Similarly, the Bengali language models BanglaBERT Bhattacharjee et al. (2022) and BanglishBERT Bhattacharjee et al. (2022) are trained on Bengali and Bengali-English corpora respectively. Code-mixed Bengali uses English tokens and hence, the pre-training on Bengali tokens does not provide any significant advantage. The lighter version of BERT, DistilBERT Sanh et al. (2019) produces comparable but slightly worse results.